Which of the following equations has only one solution?
x^2 + 4x + 4 = 0
x^2 + x = 0
x^2 - 1 = 0

Answer by TimothyLamb(4379)

---
you can solve all quadratic equations using this method:
---
x^2 + 4x + 4 = 0
---
the above quadratic equation is in standard form, with a=1, b=4, and c=4
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 4 4
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
this quadratic has ONE real root (the x-intercept), which is:
x = -2
x = -2
---
---
---
---
x^2 + x = 0
---
the above quadratic equation is in standard form, with a=1, b=1, and c=0
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 1 0
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
this quadratic has two real roots (the x-intercepts), which are:
x = 0
x = -1
---
---
---
---
x^2 - 1 = 0
---
the above quadratic equation is in standard form, with a=1, b=0, and c=-1
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 0 -1
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
this quadratic has two real roots (the x-intercepts), which are:
x = 1
x = -1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php